The equation of certain traveling waves: y(x,t)=0.0450 sin(25.12x-37.68t-0.523) where x
and y are in meters, and t in seconds. Determine the following: (a) Amplitude, (b) wave number, (c)
wavelength, (d) angular frequency, (e) frequency, (f) phase angle, (g) the wave propagation speed.
1
Expert's answer
2020-03-16T13:18:56-0400

"y(x,t)=0.0450 \sin(25.12x-37.68t-0.523)"

Comparing this with standard equation "y(x,t)=A\sin(kx-\omega t+\phi)"


Amplitude(A)=0.045 m


"k=2\pi/\lambda\\\lambda=2\pi/k"

"\lambda=2\pi/25.12=0.25"

Wavelength = 0.25


wave number = 1/ wavelength ="1\/0.25=4"


angular frequency "=\omega=37.68"

"w=2\pi v\\v= \omega/2\pi=6"

Frequency=6 Hz

Phase angle="\phi"=-0.523


Wave propagation speed="\omega/k=37.68/25.12=1.5m/s"